# KVLink: Accelerating Large Language Models via Efficient KV Cache Reuse

### v1

@@ Abstract
We describe KVLink, an approach for efficient key-value (KV) cache reuse in large language models (LLMs). In many LLM applications, different inputs can share overlapping context, such as the same retrieved document appearing in multiple queries. However, the LLMs still need to encode the entire context for each query, leading to redundant computation. In this paper, we propose a new strategy to eliminate such inefficiency, where the KV cache of each document is precomputed independently. During inference, the KV caches of retrieved documents are concatenated, allowing the model to reuse cached representations instead of recomputing them. To mitigate the performance degradation of LLMs when using KV caches computed independently for each document, KVLink introduces three key components: adjusting positional embeddings of the KV cache at inference to match the global position after concatenation, using trainable special tokens to restore self-attention across independently encoded documents, and applying mixed-data fine-tuning to enhance performance while preserving the model’s original capabilities. Experiments across 7 datasets demonstrate that KVLink improves question answering accuracy by an average of 4% over state-of-the-art methods. Furthermore, by leveraging precomputed KV caches, our approach reduces time-to-first-token by up to 90% compared to standard LLM inference, making it a scalable and efficient solution for context reuse. Code is available at https://github.com/UCSB-NLP-Chang/KVLink.
KVLink: Accelerating Large Language Models via
Efficient KV Cache Reuse
Jingbo Yang1thanks: Equal contribution. Correspondence to: Jingbo Yang $<$jingbo@ucsb.edu$>$ and Bairu Hou $<$bairu@ucsb.edu$>$  Bairu Hou1footnotemark:  Wei Wei2 Yujia Bao2 Shiyu Chang1 1UC Santa Barbara  2Center for Advanced AI, Accenture
@@ 1 Introduction
Large language models have demonstrated impressive capabilities across a broad array of applications—many of which involve processing contexts naturally divided into multiple segments. For example, in retrieval-augmented generation (RAG) (Gao et al., 2023; Wang et al., 2024a; Li et al., 2024), each retrieved document forms a distinct context chunk, while in multi-agent conversation scenarios (Wu et al., 2023; Liu et al., 2024c), outputs from different agents serve as separate segments. However, conventional autoregressive architectures require LLMs to encode the entire concatenated context as a single unit before generating a response. This approach incurs high prefilling costs for long contexts and prevents the model from separately encoding and reusing precomputed representations (i.e., key-value states) for each segment. Consider RAG: for every query, the LLM encodes a large collection of retrieved documents. When different queries share common documents, the model redundantly re-encodes these identical texts, even though their content remains unchanged.
This inefficiency motivates us to explore an alternative strategy. Instead of re-encoding the entire concatenated context for every query, we propose precomputing the key-value (KV) states for each document or text segment independently, then reusing these precomputed states during inference. By encoding each segment (e.g., each retrieved document) separately and concatenating their KV states as needed, we can eliminate redundant computations and significantly improve efficiency.
@@
However, naively encoding each document independently and concatenating their KV caches during inference can lead to performance degradation due to train-test discrepancies. Prior work (Yao et al., 2024; Sun et al., 2024; Zhang et al., 2024b) has reported up to a 35% relative decrease in accuracy on QA tasks. To overcome this challenge, we introduce KVLink, an approach designed to bridge the gap between separately encoded segments and restore self-attention across documents. KVLink introduces three key enhancements: ❶ KV cache positional re-encoding: we adjust positional embeddings during inference to ensure that the stored KV caches align correctly with the positions required for a given query. ❷ Trainable cross-segment special tokens: To effectively “link” independently encoded segments, KVLink appends a small set of trainable tokens between each segment’s precomputed KV states before concatenation. The KV representations for these tokens are computed during inference. While the documents are independently encoded into KV cache, the link tokens can attend to all the preceding tokens, which helps restore self-attention across segments while introducing only minimal computational overhead. ❸ Fine-tuning with a diverse data mixture: we fine-tune both the base LLM parameters and the newly introduced cross-segment special tokens using a diverse dataset spanning retrieval-augmented QA, summarization, and other instruction-following tasks. This process equips the model with the capability to reason across independently encoded documents while preserving its original capabilities.
@@
We validate the effectiveness and efficiency of our method through comprehensive experiments across diverse question-answering and text summarization datasets. We evaluate two backbone LLMs including Llama-3.2-1B and Llama-3.2-3B, showing that our approach consistently outperforms state-of-the-art baselines (Sun et al., 2024; Gim et al., 2024; Yao et al., 2024). For example, KVLink surpasses the best baseline by 6.6% on Natural Question (Kwiatkowski et al., 2019) and 7.3% on HotpotQA (Yang et al., 2018). Even compared to the upper-bound model (i.e., LLM using conventional full encoding for inference), our method sacrifices only minimal performance while reducing the time-to-first-token latency by up to 90% by efficiently reusing precomputed KV caches.
@@ 2.1 Problem Formulation
We aim to address the problem of reusing the key–value (KV) cache in LLMs without having to repeatedly encode the same context segments. In many applications, the same documents or context segments appear across different inference queries, yet current LLM pipelines redundantly re-encode these segments every time.
@@
Figure 1: KV cache encoding in standard and the proposed settings. In the standard approach (top), the KV cache of each document is conditioned on preceding tokens, resulting in redundant and nonreusable KV cache encodings for shared documents (e.g., $\mathrm{Doc_{b}}$). In contrast, our method (bottom) encodes documents separately, allowing KV cache reuse across queries.
@@
To address this, we consider a scenario where the KV cache for every document in the knowledge base is pre-computed in a context-free manner, as illustrated in the bottom of Figure 1. Specifically, for each document in the knowledge base, we feed only that document’s tokens into the LLM and record the resulting KV cache. At inference time, after we retrieve a set of documents for a given query, we concatenate the corresponding pre-computed KV caches. This design allows us to reuse the same cache for overlapping documents across different queries, thereby eliminating redundant computations.
However, the above approach often yields noticeable performance degradation, as LLMs are typically trained on fully concatenated sequences and each token attends to the preceding context. When we instead encode each document in a context-free manner, the model loses cross-document dependencies. Previous work also empirically demonstrates up to a 35% relative decrease in accuracy in question-answering tasks when each retrieved document is encoded into KV cache separately (Sun et al., 2024). Our method aims to enable the LLM to produce high-quality outputs when the KV cache of each document is pre-computed independently by introducing three key components, ❶ KV cache positional re-encoding. ❷ cross-document reconnection with link tokens, and ❸ mixed data fine-tuning, which will be elaborated in the following sections.
2.2 KV Cache Positional Re-encoding.
@@
More specifically, we denote the hidden state of a token at position $i$ at a particular transformer layer as $\bm{x}_{i}\in\mathbb{R}^{d}$, where $d$ is the hidden dimension. The key vector is computed as $R_{i}W_{k}\bm{x}_{i}$ and the value vector is computed as $W_{v}\bm{x}_{i}$. Here $W_{\{k,v\}}\in\mathbb{R}^{d\times d}$ represents the weight matrices that project the hidden states into the key and value spaces, and $R_{i}\in[-1,1]^{d\times d}$ is the position-dependent rotation matrix used in RoPE. Because RoPE directly encodes positional information via these rotations, the KV cache can be stored without positional embeddings, i.e., $W_{\{k,v\}}\bm{x}_{i}$. At inference time, the key-value states of all documents are concatenated, and we apply the global rotary embedding for the KV states of each token appropriate to its location in the full sequence. Note that this operation only introduces negligible time, ensuring the efficiency of our method. This is also consistent with previous methods (Yao et al., 2024; Zhang et al., 2024b), which adjust the position encoding of separately encoded KV cache during inference.
Figure 2: The attention map for all tokens. The link1 token attends only to the tokens in $\mathrm{Doc_{a}}$; link2 attends to the tokens in $\mathrm{Doc_{a}}$ and $\mathrm{Doc_{b}}$ and link2; and link3 attends to the tokens in all reused contexts and other link tokens.
Figure 3: The attention map for three link tokens and the first user input token during inference. This attention map is identical to Figure 2.
@@ 2.3 Cross-Document Reconnection with Link Tokens
2.4 Mixed Data Fine-tuning
Current LLMs are not inherently designed to handle separately encoded segments or leverage link tokens. To address this gap, we propose fine-tuning the LLM with a mixed dataset drawn from different sources, enabling the model to integrate disjoint context segments while preserving its original capabilities.
Retrieval-augmented QA.
Since RAG is the most directly relevant task, we first consider fine-tuning the model on QA tasks with retrieved documents that are separately encoded to ensure it can produce high-quality responses under this setting. More specifically, we sample questions from TriviaQA (Joshi et al., 2017) and 2WikiMQA (Ho et al., 2020). For each question, we retrieve ten relevant Wikipedia passages using Contriever (Izacard et al., 2021), each encoded separately. To obtain high-quality supervision, we prompt GPT-4 to generate reference answers, which serve as the ground truth for training.
Multi-turn conversation.
Many existing instruction-following datasets (Wang et al., 2024b) contain multi-turn conversations between the user and the assistant, which provides a natural way to train the model on disjointed contexts. We randomly convert earlier conversation turns into independently encoded KV caches, and the LLM is then trained to produce the appropriate responses in subsequent turns. This ensures the model to learn to integrate segmented contexts while maintaining its instruction-following capability.
Summarization.
Summarization serves as another useful setting where the model must aggregate information from independently encoded document chunks to generate a coherent summary. We adopt the XSum (Narayan et al., 2018) dataset. Each document is randomly split into multiple consecutive segments, and each segment is independently encoded into KV cache. The model is trained to generate the ground-truth summary, ensuring it learns to integrate information across separate segments.
Additionally, to better preserve the model’s original capabilities, we also include a standard version of TriviaQA and 2WikiMQA, where the retrieved documents are encoded as a whole rather than separately. In this setting, the model is trained to generate the ground-truth answer conditioned on the fully concatenated context. Lastly, we incorporate the T"ulu 3 dataset (Lambert et al., 2024) to preserve the instruction-following ability and a small portion of pre-training data from Fineweb (Penedo et al., 2024) to preserve the language modeling capability. More details about the dataset mixture can be found in Appendix A.1.
@@ 3.1 Experiment Setup
We evaluate our method against three existing approaches: ❶ PromptCache (Gim et al., 2024), which directly reuses the precomputed KV cache of each retrieved document or context segment for model inference. Since the original PromptCache does not handle positional encoding mismatches, we enhance it by applying our positional re-encoding method (Section2.2) to ensure the stored KV cache aligns correctly with the concatenated input. ❷ CacheBlend (Yao et al., 2024), which concatenates the precomputed KV caches of retrieved documents and then recomputes the KV cache for a small number of selected tokens within the retrieved documents. Following the original implementation, we set the recomputation ratio to 18%. ❸ BlockAttention (Sun et al., 2024), which explicitly trains the model to handle separately encoded KV caches by fine-tuning on QA tasks where retrieved documents are encoded independently. For all baselines and our method, we use Llama-3.2-1B-Instruct and Llama-3.2-3B-Instruct (Meta AI, 2024) as the backbone models. Since BlockAttention and our method require fine-tuning, we ensure fair comparison by using the same data mixture and training hyperparameters for both. Further implementation details for each baseline are provided in Appendix A.3.
@@
We adopt Llama-3.2-1B-Instruct and Llama-3.2-3B-Instruct as the backbone models, fine-tuning them for 6,000 steps using a global batch size of 64 across 8 A100 GPUs. Further details on data preprocessing, dataset mixture, and training configurations are provided in Appendix A.2. For our method, we train three versions, each appending 0, 1, or 5 link tokens to every document or context segment, denoted as KVLink0, KVLink1, and KVLink5, respectively.
@@
We evaluate the effectiveness of our method in three dimensions: ❶ performance with separately encoded KV cache, ❷ the inference efficiency, and ❸ the general capability preservation (e.g., math reasoning and instruction-following ability).
@@
To evaluate the model performance when using separately encoded documents or context segments, we focus on retrieval-augmented question answering tasks including NaturalQuestions Kwiatkowski et al. (2019), 2WikiMQA Ho et al. (2020), TriviaQA Joshi et al. (2017), HotpotQA Yang et al. (2018), and MuSiQue Trivedi et al. (2022). For NaturalQuestions, we adopt the evaluation protocol from Liu et al. (2024a), where the LLM is given a question along with a set of 10 retrieved documents. The document that contains the correct answer is systematically placed at each of the 10 possible positions across separate evaluation runs. The final accuracy is reported as the average over these 10 evaluations. For 2WikiMQA, HotpotQA, and MuSiQue, we utilize the originally provided retrieved documents for evaluation. For TriviaQA we retrieve 10 documents using Contriever (Izacard et al., 2021) following the setting in BlockAttention. In all cases, documents are encoded separately into KV cache.
Additionally, following CacheBlend, we evaluate on two text summarization datasets including MultiNews (Fabbri et al., 2019) and Samsum (Gliwa et al., 2019). The LLM is given several in-context examples per instance, each separately encoded into KV cache. We report the RougeL score (Lin, 2004) as the metric. Notably, our evaluation datasets cover all those used by our baselines, BlockAttention and CacheBlend, ensuring a comprehensive and fair comparison.
@@
Finally, to ensure our method does not degrade the model’s original capabilities, we evaluate it on a range of reasoning and instruction-following benchmarks, including IFEval (Zhou et al., 2023), GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2020), ARC-Challenge (Clark et al., 2018), ARC-Easy (Clark et al., 2018), PiQA (Bisk et al., 2020), SciQ (Welbl et al., 2017), Winogrande (Sakaguchi et al., 2021), and HellaSwag (Zellers et al., 2019). We report the accuracy on each dataset. For IFEval, we report both strict prompt-level accuracy (IFEval-P) and instruction-level accuracy (IFEval-I). More details of the evaluation configuration are available in Appendix A.2.

### v2

@@ Abstract
We describe KVLink, an approach for efficient key-value (KV) cache reuse in large language models (LLMs). In many LLM applications, different inputs can share overlapping context, such as the same retrieved document appearing in multiple queries. However, the LLMs still need to encode the entire context for each query, leading to redundant computation. In this paper, we investigate a new strategy to eliminate such inefficiency, where the KV cache of each document is precomputed independently. During inference, the KV caches of retrieved documents are concatenated, allowing the model to reuse cached representations instead of recomputing them. To mitigate the performance degradation when using KV caches computed independently for each document, KVLink introduces two key techniques: adjusting positional embeddings of the KV cache at inference to match the global position after concatenation, and using trainable special tokens to restore self-attention across independently encoded documents. Experiments across 7 datasets demonstrate that KVLink improves question answering accuracy by an average of 4% over state-of-the-art methods. Furthermore, by leveraging precomputed KV caches, our approach reduces time-to-first-token by up to 96% compared to standard LLM inference, making it a scalable and efficient solution for context reuse. Additionally, KVLink can be combined with KV cache compression to further save cache loading and storage overhead while outperforming the baselines. Code is available in the supplementary material.
@@ 1 Introduction
Large language models have demonstrated impressive capabilities across a broad array of applications—many of which involve processing contexts naturally divided into multiple segments. For example, in retrieval-augmented generation (RAG) [1, 2, 3], each retrieved document forms a distinct context chunk, while in multi-agent conversation scenarios [4, 5], outputs from different agents serve as separate segments. However, conventional architectures require LLMs to encode the entire concatenated context as a single unit before generating a response. This approach incurs high prefilling costs for long contexts and prevents the model from separately encoding and reusing precomputed representations (i.e., key-value states) for each segment. Consider RAG: for every query, the LLM encodes a large collection of retrieved documents. When different queries share common documents, the model redundantly re-encodes these identical texts, even though their content remains unchanged.
@@
This inefficiency motivates us to explore an alternative strategy. Instead of re-encoding the entire concatenated context for every query, we propose precomputing the key-value (KV) states for each document or text segment independently, then reusing these precomputed states during inference. By encoding each segment (e.g., each retrieved document) separately and concatenating their KV states as needed, we can eliminate redundant computations and significantly improve efficiency. However, naively encoding each document independently and concatenating their KV caches during inference can lead to performance degradation due to train-test discrepancies. Prior work [6, 7, 8] has reported up to a 35% relative decrease in accuracy on QA tasks, which results from the discrepancy of position embeddings and missing cross attention between retrieved documents. To overcome this challenge, we introduce KVLink, an approach designed to bridge the gap between separately encoded segments and restore self-attention across documents. KVLink introduces two key enhancements: ❶ KV cache positional re-encoding. We adjust positional embeddings during inference to ensure that the stored KV caches align correctly with the positions required for a given query. ❷ Trainable cross-segment special tokens. To effectively “link” independently encoded segments, KVLink appends a small set of trainable tokens between each segment’s precomputed KV states before concatenation. The KV representations for these tokens are computed during inference. While the documents are independently encoded into KV cache, the link tokens can attend to all the preceding tokens, which helps restore self-attention across segments while introducing only minimal computational overhead.
Figure 1: Standard approach (top) encodes the KV cache of each document conditioned on preceding tokens, resulting in redundant and nonreusable KV cache encoding for shared documents (e.g., $\mathrm{Doc_{b}}$). In contrast, our setting (bottom) encodes documents separately, allowing KV cache reuse across queries.
@@
We validate the effectiveness and efficiency of our method through comprehensive experiments across diverse question-answering and text summarization datasets. We evaluate three backbone LLMs with different model scales, including Llama-3.2-1B, Llama-3.2-3B, and Llama-3.1-8B, showing that our approach consistently outperforms state-of-the-art baselines [6, 7, 9]. For example, KVLink surpasses the best baseline by 6.6% on Natural Question [10] and 7.3% on HotpotQA [11]. Even compared to the upper-bound model (i.e., LLM using conventional full encoding for inference), our method sacrifices only minimal performance while reducing the time-to-first-token latency by up to 96% by efficiently reusing precomputed KV caches.
Additionally, we investigate the integration of KVLink with KV cache compression techniques, as storing precomputed KV caches can incur substantial storage overhead. By incorporating methods such as LLMLingua [12] and AnLLMs [13], KVLink consistently outperforms baseline approaches. These results demonstrate its strong potential for practical, large-scale deployment.
@@ 2.1 Problem Formulation
We aim to address the problem of reusing the key-value (KV) cache in LLMs without having to repeatedly encode the same context segments. In many applications, the same documents or context segments appear across different inference queries, yet current LLM pipelines redundantly re-encode these segments every time.
@@
To address this, we consider a scenario where the KV cache for every document in the knowledge base is pre-computed in a context-free manner, as illustrated in the bottom of Figure 1. Specifically, for each document in the knowledge base, we feed only that document’s tokens into the LLM and record the resulting KV cache. At inference time, after we retrieve a set of documents for a given query, we concatenate their pre-computed KV caches. This design allows us to reuse the same cache for overlapping documents across different queries, thereby eliminating redundant computations.
However, the above approach often yields noticeable performance degradation, as LLMs are typically trained on fully concatenated sequences and each token attends to the preceding context. When we instead encode each document in a context-free manner, the model loses cross-document dependencies. Previous work also empirically demonstrates up to a 35% relative decrease in accuracy in question-answering tasks when each retrieved document is encoded into KV cache separately [7]. Our method aims to enable the LLM to produce high-quality outputs when the KV cache of each document is pre-computed independently by introducing two key components, ❶ KV cache positional re-encoding and ❷ cross-document reconnection with link tokens.
2.2 KV Cache Positional Re-encoding
@@
More specifically, we denote the hidden state of a token at position $i$ at a particular transformer layer as $\bm{x}_{i}\in\mathbb{R}^{d}$, where $d$ is the hidden dimension. The key vector is computed as $R_{i}W_{k}\bm{x}_{i}$ and the value vector is computed as $W_{v}\bm{x}_{i}$. Here $W_{\{k,v\}}\in\mathbb{R}^{d\times d}$ represents the weight matrices that project the hidden states into the key and value spaces, and $R_{i}\in[-1,1]^{d\times d}$ is the position-dependent rotation matrix used in RoPE. Because RoPE directly encodes positional information via these rotations, the KV cache can be stored without positional embeddings, i.e., $W_{\{k,v\}}\bm{x}_{i}$. At inference time, the key-value states of all documents are concatenated, and we apply the global rotary embedding for the KV states of each token appropriate to its location in the full sequence. This operation only introduces negligible time, ensuring the efficiency of our method. This is also consistent with previous methods [6, 8], which adjust the position encoding of separately encoded KV cache during inference.
Figure 2: Left: the attention map for all tokens. The link1 token attends only to the tokens in $\mathrm{Doc_{a}}$; link2 attends to the tokens in $\mathrm{Doc_{a}}$ and $\mathrm{Doc_{b}}$ and link2; and link3 attends to the tokens in all reused contexts and other link tokens. Right: the attention map for three link tokens and the first user input token during inference. These two attention maps are identical.
@@ 2.3 Cross-Document Reconnection with Link Tokens
2.4 Compressed KV Cache Linking
Although reusing pre-computed KV caches can greatly accelerate inference, it may introduce considerable overhead in storing these caches in a database and loading them onto the GPU. This overhead presents a major challenge for integrating cache reuse into real-world RAG deployments. For example, a 1,000-token document stored as UTF-8 text occupies only about 5KB, whereas the corresponding Llama3-8B KV cache requires roughly 131MB. To address this issue, we further explore combining our method with KV cache compression techniques. Specifically, we consider the following choices: ❶ LLMLingua [12], which compresses documents by token dropping; and ❷ AnLLMs [13], which fine-tunes the LLM to compress each sentence in the document into an anchor token with special attention masks.
When using AnLLMs, we empirically observe a significant drop in LLM performance on question-answering benchmarks. To address this, we introduce the following modifications aimed at improving effectiveness. First, instead of compressing each sentence into a single anchor token and storing its KV cache, we divide each document into multiple consecutive chunks of fixed length and compress each chunk into multiple anchor tokens. Second, we modify the attention masks so that each anchor token attends only to tokens within its own chunk and to preceding anchor tokens. More specifically, let $\bm{x}$ represent a training example from the pretraining dataset, consisting of a sequence of tokens. We split the first $N$ tokens into $N/s$ chunks of fixed size $s$:
$\bm{x}^{(n)}=x_{(n-1)s+1},\dots,x_{ns},\,\,n=1,\dots,N,$
where $s$ denotes the chunk size. In our experiments, we set $s=100$ by default. We then append several anchor tokens to each chunk. The LLM is trained to perform next-token prediction on the remaining tokens beyond position $N$, which are restricted to attend only to the anchor tokens. This setup encourages the model to compress the input into a small number of anchor tokens, allowing us to store only the KV cache of those tokens and thereby significantly reduce storage overhead. Also, by increasing the number of anchor tokens, we can better maintain the model performance given compressed KV cache. To apply this method, we first pretrain the LLM using this objective on pretraining data. We then fine-tune the model on QA datasets using KVLink. Additional implementation details are provided in Appendix A.6.
@@ 3.1 Experiment Setup
We evaluate our method against three existing approaches: ❶ PromptCache [9], which directly reuses the precomputed KV cache of each retrieved document or context segment for model inference. Since the original PromptCache does not handle positional encoding mismatches, we enhance it by applying our positional re-encoding method (Section2.2) to ensure the stored KV cache aligns correctly with the concatenated input. ❷ CacheBlend [6], which concatenates the precomputed KV caches of retrieved documents and then recomputes the KV cache for a small number of selected tokens within the retrieved documents. Following the original implementation, we set the recomputation ratio to 18%. ❸ BlockAttention [7], which explicitly trains the model to handle separately encoded KV caches by fine-tuning on QA tasks where retrieved documents are encoded independently. For all baselines and our method, we use Llama-3.2-1B-Instruct and Llama-3.2-3B-Instruct [15] as the backbone models. Since BlockAttention and our method require fine-tuning, we ensure fair comparison by using the same data mixture and training hyperparameters for both. Further implementation details for each baseline are provided in Appendix A.3.
@@
We adopt Llama-3.2-1B-Instruct, Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct as the backbone models, fine-tuning them for 6,000 steps using a global batch size of 64 across 8$\times$×H100 GPUs. We construct the training dataset by mixing the training sets of 2WikiMQA , TriviaQA , pretraining data from FineWeb [16], and Tülu 3 [17]. Further details on data preprocessing, dataset mixture, and training configurations are provided in Appendix A.1 and  A.2. For our method, we train three versions, each appending 0, 1, or 5 link tokens to every document or context segment, denoted as KVLink0, KVLink1, and KVLink5, respectively.
@@
We evaluate the effectiveness of our method in three dimensions: ❶ performance with separately encoded KV cache, ❷ the inference efficiency, ❸ the general capability preservation (e.g., math reasoning and instruction-following ability), and ❹ performance with compressed KV cache.
@@
Llama-3.1-8B
Original Llama 71.3% 77.2% 76.2% 77.6% 49.1% 0.204 0.348
Finetuned Upperbound 75.1% 78.5% 77.9% 78.2% 46.5% 0.166 0.353
PromptCache 28.9% 42.2% 62.6% 36.5% 6.7% 0.203 0.329
CacheBlend 55.2% 45.9% 68.8% 40.8% 6.0% 0.207 0.320
BlockAttention 70.8% 73.6% 76.6% 72.2% 38.7% 0.166 0.342
KVLink 0 71.0% 73.3% 76.4% 72.7% 39.9% 0.163 0.342
KVLink 1 72.4% 74.7% 76.6% 73.6% 38.9% 0.170 0.340
KVLink 5 72.5% 73.8% 76.6% 73.8% 40.8% 0.168 0.345
To evaluate the model performance when using separately encoded documents or context segments, we focus on retrieval-augmented question answering tasks including NaturalQuestions [10], 2WikiMQA [18], TriviaQA [19], HotpotQA [11], and MuSiQue [20]. For NaturalQuestions, we adopt the evaluation protocol from Liu et al. [21], where the LLM is given a question along with a set of 10 retrieved documents. The document that contains the correct answer is systematically placed at each of the 10 possible positions across separate evaluation runs. The final accuracy is reported as the average over these 10 evaluations. For 2WikiMQA, HotpotQA, and MuSiQue, we utilize the originally provided retrieved documents for evaluation. For TriviaQA we retrieve 10 documents using Contriever [22] following the setting in BlockAttention. In all cases, documents are encoded separately into KV cache.
Additionally, following CacheBlend, we evaluate on two text summarization datasets including MultiNews [23] and Samsum [24]. The LLM is given several in-context examples per instance, each separately encoded into KV cache. We report the RougeL score [25] as the metric. Notably, our evaluation datasets cover all those used by our baselines, BlockAttention and CacheBlend, ensuring a comprehensive and fair comparison.
@@
To ensure our method does not degrade the model’s original capabilities, we evaluate it on a range of reasoning and instruction-following benchmarks, including IFEval [26], GSM8K [27], MMLU [28], ARC-Challenge [29], ARC-Easy [29], PiQA [30], SciQ [31], Winogrande [32], and HellaSwag [33]. We report the accuracy on each dataset. For IFEval, we report both strict prompt-level accuracy (IFEval-P) and instruction-level accuracy (IFEval-I). More details of the evaluation configuration are available in Appendix A.2.
Finally, to further reduce cache-storage and loading overhead, and thereby make cache reuse practical, we evaluate reuse with compressed KV caches. We report QA accuracy under various compression methods and assess the effectiveness of KVLink when operating on the compressed caches. The only difference from the first experiment is that we use compressed, rather than original, KV cache.
3.2 Main Results
We highlight the key observations below. First, KVLink consistently outperforms all baselines across all datasets, demonstrating its strong effectiveness. In most QA tasks, our approach surpasses the best baseline, BlockAttention, with up to 5% higher accuracy. The only exceptions are the 2WikiMQA and TriviaQA datasets, whose training sets are included in our training data mixture. On all other evaluation datasets, our method consistently outperforms the baselines, demonstrating superior out-of-domain generalization performance. For other baselines, the gap is more significant. The performance gap is even larger when compared to other baselines. Notably, our method not only outperforms the original Llama models without fine-tuning but also achieves accuracy close to fine-tuned Llama, which is evaluated with a fully concatenated context.
3.3 Inference Efficiency Evaluation
A key objective of KVLink is to reduce the computational overhead incurred by repeatedly encoding long contexts. To demonstrate these efficiency gains, we measure the Time-to-First-Token (TTFT) when reusing the precomputed KV cache for ten documents of varying lengths. For a realistic assessment, we include the overhead of loading the precomputed KV caches onto the GPU. Each measurement is averaged over 100 runs, with an initial 10 warm-up trials to eliminate memory allocation overhead, following the setup in  [34].
Figure 3: Inference speed comparison with ten reused contexts of varying lengths. Both KVLink1 and KVLink5 show considerably lower Time-to-First-Token (TTFT) than standard decoding as context size grows.
Figure 3 compares our methods, KVLink1 and KVLink5, against standard decoding, which encodes the entire context for each query. The key observations are as follows. First, reusing the precomputed KV cache significantly improves LLM inference efficiency. To leverage the precomputed KV cache, we perform three main operations: (1) loading KV caches from CPU to GPU, (2) applying new positional encoding, and (3) encoding the link tokens for each document. All of these operations introduce only negligible latency, reducing TTFT by 85%–96% compared to standard decoding.
Second, across all context sizes, both KVLink1 and KVLink5 consistently achieve substantially lower TTFT relative to standard decoding. As the reused context length increases, the TTFT gap further widens, highlighting the scalability of our approach. Notably, when the context length reaches 5,000 tokens, KVLink decreases inference latency by 96%. These results confirm that KVLink effectively mitigates the computational cost of large-scale context encoding and is well-suited for scenarios requiring fast response times.
3.4 General Capability Preservation
While our primary objective is to enable efficient KV cache reuse, we also verify whether these modifications maintain the model’s general reasoning and instruction-following capabilities. Table 2 presents the results for three model variants: Llama-3.2-1B, Llama-3.2-3B and Llama-3.1-8B. We compare the original models, their fine-tuned counterparts, and our method with different numbers of link tokens (KVLink0, KVLink1, and KVLink5).